OpenStudy (anonymous):

If there is 1 child, there can be 0 interactions. If there are 2 children, there can be 3 interactions. If there are 3 children, there can be 9 interactions. If there are 4 children, there can be 18 interactions. How many interactions will there be for 9 children? 135 84 93 108

OpenStudy (rsadhvika):

observe the pattern :- \(a_{n} = a_{n-1} + 3(n-1)\)

OpenStudy (rsadhvika):

find the interactions upto 9 childern using above recursive relation

OpenStudy (anonymous):

what do i plug in for the variables?

OpenStudy (rsadhvika):

lets find number of interactions for 5 children :- \(a_{5} = a_{5-1} + 3(5-1) \) \(a_{5} = a_{4} + 3(4) \) \(a_{5} = 18 + 12 \) \(a_{5} = 30 \)

OpenStudy (rsadhvika):

similarly, find a6, a7, a8 and a9

OpenStudy (anonymous):

where did the 18 come from?

OpenStudy (rsadhvika):

good q, look at the orignial quesiton

OpenStudy (rsadhvika):

If there are 4 children, there can be 18 interactions.

OpenStudy (anonymous):

mhmm